OpenStudy (anonymous):

Use a coterminal angle to find the exact value of the expression. Do not use a calculator. Seleccione una respuesta. a. b. c. d.

OpenStudy (anonymous):

are you spanish? and where is the question and options

OpenStudy (anonymous):

\[\cos405\]

OpenStudy (anonymous):

cos405=cos(360+45)=cos45=1/root(2)

OpenStudy (anonymous):

to choose one answer

OpenStudy (anonymous):

No tengo idea de ángulo coterminal ....... lo siento

OpenStudy (anonymous):

but the answer i found is.707

OpenStudy (ash2326):

co405=cos(360+45) cos (A+B)=cosAcosB-sinAsinB cos(360+45)=cos(360)cos(45)-sin(360)sin(45) cos360=1 sin360=0 sin45=cos45=1/sqrt(2) so cos405=1/sqrt(2) or 0.707

OpenStudy (ash2326):

\[1/\sqrt(2)\]=1/1.414=0.707

OpenStudy (anonymous):

but ash2326 , in options there is no .707

OpenStudy (ash2326):

you wrote before that answer given is 0.707

OpenStudy (anonymous):

Use a coterminal angle to find the exact value of the expression. Do not use a calculator. \[\cos 405\]° Choose the rifht answer: